# PuzzleWorld: A Benchmark for Multimodal, Open-Ended Reasoning in Puzzlehunts

### v1

@@ 4.2 Data annotation
To facilitate the development of reasoning capabilities in AI systems, we designed a comprehensive annotation structure for our puzzle dataset. Each puzzle in the corpus is represented by a standardized set of metadata and accompanying visual assets, enabling systematic analysis of reasoning patterns and solution pathways. To prevent ambiguity, we discard puzzles that have incomplete solutions, multiple ground truth answers, or require physical activity to solve the puzzles. This leaves us with 667 annotated puzzles.
Figure 4: Illustration of metadata schema: All puzzles are annotated with their accompanying metadata, which includes the puzzle title, flavor text, difficulty, final answer, reasoning steps, input modalities, reasoning skills, and the PuzzledPint link to the puzzle.
@@ 4.2.1 Metadata schema
Each puzzle is annotated using a structured JSON schema comprising several key fields: a title that serves as a descriptive identifier; optional flavor text that provides narrative context; a difficulty label (easy, medium, or hard) indicating the puzzle’s complexity; a solution representing the canonical answer; a reasoning field outlining an ordered sequence of steps leading to the solution; a modality tag specifying the types of input involved; a list of skills capturing the cognitive abilities required for solving; and a source field attributing the data origin. Figure 4 illustrates an annotation schema example.
@@ 4.2.2 Reasoning annotation
A key contribution of our annotation methodology is the decomposition of puzzle-solving into discrete reasoning steps. Each step is formalized as a tuple $\langle e,f\rangle$ where $e$ represents the textual explanation and $f$ denotes an optional figure illustrating the reasoning process. To ensure consistency across annotations, we loosely require each step to begin with an atomic cognitive operation, such as pattern discovery or sketching, followed by the intermediate outcome of that operation. This structured annotation enables fine-grained analysis of a model’s reasoning trajectory.
4.3 Data verification
To ensure annotation quality and integrity, we implemented a rigorous two-stage verification protocol. First, we used GPT-4o to automatically flag each puzzle annotation for correctness and reasoning coherence. This automated screening specifically identified reasoning steps exhibiting ambiguity or logical discontinuities that might impede systematic analysis, which has flagged 12.11% of the puzzle dataset. Subsequently, two human verifiers independently reviewed all flagged annotations, applying corrections where necessary. This verification process resulted in modifications to 10.93% of the initially annotated puzzles.
As an additional quality assurance measure, we conducted manual verification of a randomly selected subset comprising 5% of the total dataset. This evaluation revealed that 96.5% of the verified annotations are marked as correct by the verifiers, demonstrating the high reliability of our annotation methodology.
@@ 4.4 Dataset statistics
We summarize key statistics in Figure 3 (right). The average number of reasoning steps is above 5, and the average word count per reasoning step is above 20, demonstrating the complexity of the reasoning traces. Additionally, 12.3% of the steps have a visual intermediate output, highlighting the importance of sketching and spatial reasoning to solve puzzles. The correlation between puzzle difficulty and # of reasoning steps is 0.24. While we expect difficulty and # of reasoning steps to be positively correlated, the magnitude of the correlation is relatively low, as the difficulty of the puzzles also stems from their open-ended nature. Figure 5 shows the distribution of puzzles by modalities, reasoning skills, number of reasoning steps, and difficulty.
Figure 5: PuzzleWorld dataset statistics. Distributions of modalities and reasoning skills are balanced across different puzzles. While the majority of puzzles are classified as medium difficulty, there are a significant number of easy and hard puzzles. The number of reasoning steps follows a long-tail distribution, with many puzzle solutions requiring more than 5 steps and some hard puzzles requiring up to 30 steps of reasoning.
4.5 Checking for dataset contamination
To assess the possibility of data contamination, we test whether GPT-o3 [28] has memorized any of the puzzles in our dataset. Specifically, inspired by prior work [33, 7], we prompt the model to reconstruct the flavor text for 40 randomly sampled puzzles out of the 84 that were answered correctly. We then use GPT-4o [2] to automatically evaluate the similarity between the reconstructed and original flavor texts. We find a reconstruction accuracy of 0%, suggesting little to no evidence of data leakage. Furthermore, since Puzzled Pint [29] publishes new puzzles on a monthly basis, our dataset can be continuously updated to mitigate the risk of model overfitting on released content.
@@ 5.1 Experimental setup
We evaluate state-of-the-art closed-source and open-source multimodal large language models on PuzzleWorld. We select GPT-o3 and GPT-4o as closed-source candidates, along with open-source models Qwen QVQ[35], InternVL3[51], and Kimi VL A3B[34]. Following the protocol in Wang et al. [37], we employ a standardized prompting strategy: the model is first introduced to the puzzle-solving task with general hints, followed by the puzzle content, including images and transcribed flavor text. See Appendix C for the full prompt.
@@ 5.1.1 Automatic evaluation metrics
Beyond the models’ final answer accuracy, we additionally evaluate the models’ solution stepwise accuracy by comparing their solution reasoning with the annotated ground truth reasoning steps. Since puzzles can have multiple solution pathways, we define the stepwise accuracy score of a candidate solution to be the last annotated reasoning step it successfully identified and executed out of all the reasoning steps.
To automatically evaluate candidate responses from multiple models, we implement an LLM-as-a-judge [49] with GPT-4o to determine the score of each candidate solution. The LLM judge takes in the grading prompt, the candidate response, and the reference reasoning steps from our dataset. For each reasoning step in the reference solution, the LLM judge then determines if the step is met by the candidate response. To evaluate the reliability of the LLM judge, we compared its stepwise accuracy scores on 20 randomly selected puzzles against human-annotated evaluations. The LLM judge achieved a Pearson correlation of $r=0.829$ ($p=6.3\times 10^{-6}$) and a mean absolute error (MAE) of $0.083$ with respect to human scores, indicating strong alignment with human judgment.
@@ 5.2.1 Overall performance of frontier models
Table 2: Model performance. Accuracy (Acc) and stepwise accuracy (Step) are reported overall and per modality: text, visual, and structured. Today’s state-of-the-art models struggle significantly on PuzzleWorld- most achieve only 1-2% final answer accuracy, with the best model (GPT-o3, an expensive and slow reasoning model) solving only 14% of puzzles and reaching 40% stepwise accuracy.
@@
We report the performance of both open-source and closed-source models in Table 2. All models exhibit extremely low final answer accuracy on PuzzleWorld, with most achieving close to 1-2%. GPT-o3 attains the highest overall accuracy at 14.22%, while the best-performing open-source model, QVQ-72B-Preview, reaches just 1.36%. Although these uniformly low scores underscore the difficulty of our benchmark, they offer limited insight into the models’ reasoning capabilities.
To address this, we introduce stepwise evaluation metrics that provide a more nuanced view of intermediate reasoning performance. These metrics reveal that models with poor final answer accuracy, such as InternVL3, can still demonstrate good intermediate reasoning, achieving up to 15.49% stepwise accuracy. Similarly, while QVQ-72B-Preview lags behind GPT-4o in final answer accuracy, it outperforms it in stepwise accuracy (30.2%), reflecting more coherent reasoning trajectories despite not reaching the correct final output. These two metrics enable PuzzleWorld to remain highly challenging while offering detailed diagnostics for model evaluation and development.
In terms of input modalities, models generally perform best on text-based puzzles, with significantly lower accuracy on puzzles involving unstructured visual inputs. Interestingly, some models’ performance improves on structured visual puzzles, such as crosswords and logic grids, where the spatial format is regular and constrained. For example, GPT-o3 performs reliably on simple crosswords that have clearly defined conventions. In contrast, puzzles involving free-form or unconventional visual layouts remain particularly difficult, with models often achieving less than half their text puzzle accuracy on these inputs. These trends highlight persistent weaknesses in visual grounding and spatial reasoning across current models.
Table 3: Fine-tuning model accuracy on PuzzleWorld. Stepwise accuracy improves significantly when fine-tuning Intern-VL3 on reasoning traces, while final answer accuracy remains unchanged.
@@
5.3 Improving reasoning on PuzzleWorld with fine-tuning
To explore whether PuzzleWorld can support model improvement, we fine-tuned a smaller 8B Intern-VL3 model on 80% of the dataset using annotated reasoning traces, and evaluated performance on the 20% held-out test set. As a control, we fine-tuned the same model using only the final answers, without access to intermediate reasoning traces. Full fine-tuning details are provided in Appendix C.2.
Our results highlight the value of PuzzleWorld’s reasoning annotations. Fine-tuning on full reasoning traces more than doubles the model’s stepwise accuracy—from 4.78% (base model) to 11.00%. In contrast, fine-tuning on final answers alone significantly impairs performance, reducing stepwise accuracy to 2.96% and driving final answer accuracy to zero. This suggests that answer-only training can overfit to the answer only and disrupt coherent reasoning behavior.
Despite these improvements, the final answer accuracy remains unchanged at 0.76%, underscoring both the difficulty of our benchmark and the limitations of standard fine-tuning approaches in addressing the reasoning challenges in PuzzleWorld. These findings demonstrate the importance of our detailed reasoning annotations and more advanced training paradigms. We now present a detailed error analysis to identify models’ specific failure modes on PuzzleWorld that can inform future directions for multimodal reasoning.
@@ 5.4 Detailed error analysis
Figure 6: Example puzzle errors. Left: (myopic reasoning) The model outputs an incorrect plan to solve the puzzle and is unable to successfully backtrack when it hits a dead end. Middle: (language bottleneck/lack of visual understanding) The model misinterprets the visual contents of the puzzle due to inherent limitations in language. Right: (sketching errors) While the model may produce a plausible plan, it fails to execute the right visual sketching steps to obtain correct intermediate outputs. High-resolution puzzle images are in Appendix B.
We highlight the main sources of errors by the best reasoning multimodal LLMs on PuzzleWorld, focusing on GPT-o3, which achieved best results (14% final answer accuracy). See Figure 6 for example errors from each category.
@@
Figure 7: Stepwise accuracy distribution of GPT-o3. GPT-o3 receives stepwise accuracy of 0 for most puzzles, indicating the model frequently fails to identify even the first step of the correct reasoning trace. This highlights GPT-o3’s tendency toward myopic reasoning and its inability to recover or backtrack once committed to an incorrect path.
Strong models exhibit myopic reasoning.
@@
Despite strong performance on conventional reasoning benchmarks, frontier multimodal models often exhibit myopic commitment in their chain-of-thought behavior. Rather than flexibly exploring alternatives or revisiting prior steps, models tend to fixate on an early hypothesis–often triggered by surface-level cues–resulting in reasoning trajectories that are locally coherent but globally misaligned with the puzzle content. For example, in Figure 6, solving the puzzle requires interpreting musical notes using a mix of binary, Morse code, and flag semaphores. Instead, GPT-o3 identifies a Morse code reference early on and rigidly adheres to it–even as contradictions arise–demonstrating a lack of backtracking and global verification.
@@
To further examine this behavior, we analyze the stepwise accuracy distribution of GPT-o3 (Figure 7). We find that, on most puzzles, the model receive a score of 0, meaning the model often fails to correctly identify even the first step of the annotated solution. Once committed to an incorrect path, the model rarely recovers, highlighting the brittle nature of its reasoning and its lack of mechanisms for self-correction or consistency checking.
Multimodal models are bottlenecked by the limitations of language.
Modern multimodal models rely heavily on language-based strategies, such as long-form chain-of-thought and code generation, for reasoning. However, we find that this dependence becomes a bottleneck in puzzles with complex visual structure. In Figure 6, the puzzle is composed of four interlocking, ring-shaped loops arranged in a clover-like pattern. This layout is intuitive to navigate visually, as the spatial flow naturally guides solvers through each ring. However, the same structure is difficult to represent linearly in text.
@@
Figure 8: Example failure case where GPT-o3 fails to convert a complex structured puzzle into text.
@@
While GPT-o3 correctly solves the word clues within each loop, it fails to capture the global layout when converting the puzzle into text, as shown in Figure 8. This ultimately leads the model to derive an incorrect final answer. This example highlights a broader limitation: when faced with highly structured multimodal inputs, models that default to textual reasoning often lose critical spatial information, revealing the inherent mismatch between visual intuition and language-centric inference.
@@
While frontier models have made notable progress in logical deduction and arithmetic reasoning, they consistently underperform on spatial tasks that require sketching, drawing, and manipulating visual structure, such as decoding based on spatial arrangements or tracing paths through grids and mazes. In Figure 6, for instance, the model correctly solves the individual clues in a grid-based puzzle but fails to trace the intended path, resulting in an incorrect final answer. Humans naturally rely on sketching or mental imagery to reason through such spatial challenges, using external or internal visualizations to keep track of evolving structure. The absence of such capabilities in current models reveals a critical gap: without the ability to sketch and update a persistent visual representation during reasoning, models are prone to failure in tasks that depend on spatial coherence.
Figure 9: Reasoning skills of failed steps. We annotated the reasoning steps responsible for 30 puzzle errors with their corresponding reasoning skills.
To better understand the role of sketching in model performance, we manually analyzed 30 puzzles where GPT-o3 produced incorrect answers. For each failure case, we annotated the reasoning step responsible for the error with its corresponding reasoning skill category from the dataset. As shown in Figure 9, we found that 53.33% of these bottleneck steps involved spatial reasoning or sketching-related capabilities. This highlights a significant gap in current models’ ability to maintain and manipulate visual structure during inference. Incorporating sketch-like visual memory and spatial reasoning mechanisms [43, 16, 6] may offer a promising direction for building more robust and spatially grounded multimodal reasoning AI.

### v2

@@ 4.2 Data annotation
To facilitate AI’s reasoning capabilities, we designed a comprehensive annotation structure for PuzzleWorld. Each puzzle is represented by a standardized metadata and visual assets. To prevent ambiguity, we discard puzzles that have incomplete solutions, multiple ground truth answers, or require physical activity to solve the puzzles. This leaves us with 667 annotated puzzles.
Figure 4: Illustration of metadata schema: All puzzles are annotated with accompanying metadata, which includes the title, flavor text, difficulty, final answer, reasoning steps, input modalities, reasoning skills, and the link to the puzzle.
@@ 4.2.1 Metadata schema
Each puzzle is annotated using a JSON schema comprising several fields: a descriptive title; flavor text providing narrative context; a difficulty label (easy, medium, or hard); solution representing the canonical answer; a reasoning field of an ordered sequence of steps leading to the solution; a modality tag specifying the input types involved; a list of skills capturing the cognitive abilities required for solving; and a source field attributing the data origin. Figure 4 illustrates an example annotation.
@@ 4.2.2 Reasoning annotation
A key contribution of our annotations is the decomposition of puzzle-solving into reasoning steps. Each step is formalized as a tuple $\langle e,f\rangle$ where $e$ represents the text explanation and $f$ denotes an optional figure illustrating the reasoning. To ensure consistency, we loosely require each step to begin with an atomic operation, such as pattern discovery or sketching, followed by the operation’s intermediate output. This annotation structure enables fine-grained reasoning trajectory analysis on AI solutions.
4.3 Verification of Annotations and Data Contamination
To ensure annotation quality and integrity, we implemented a two-stage verification protocol. First, we used GPT‑4o to flag each puzzle annotation for correctness and reasoning coherence. This automated screening identified reasoning steps exhibiting ambiguity or logical discontinuities that might impede systematic analysis, which has flagged 12.11% of the dataset. Subsequently, two human verifiers independently reviewed all flagged annotations, applying corrections where necessary. This verification process resulted in modifications to 10.93% of the initially annotated puzzles. As an additional quality assurance measure, we conducted manual verification of a random subset comprising 5% of the dataset. In this evaluation, 96.5% of the verified annotations are marked as correct by the verifiers, demonstrating the high reliability of our annotation methodology. Finally, we verify whether frontier models has memorized any of the puzzles in PuzzleWorld. We describe our procedure in E.1, where we find no evidence of data contamination.
@@ 4.4 Dataset statistics
We summarize key statistics in Figure 3 (right). The average number of reasoning steps is above 5, and the average word count per reasoning step is above 20, demonstrating the complexity of the reasoning traces. Additionally, 12.3% of the steps have a visual intermediate output, highlighting the importance of sketching and spatial reasoning to solve puzzles. The correlation between puzzle difficulty and number of reasoning steps is 0.24. While we expect difficulty and number of reasoning steps to be positively correlated, the magnitude of the correlation is relatively low, as the difficulty of the puzzles also stems from their open-ended nature. Figure 5 shows the distribution of puzzles by modalities, reasoning skills, number of reasoning steps, and difficulty.
Figure 5: PuzzleWorld dataset statistics. Distributions of modalities and reasoning skills are balanced. While the majority of puzzles are of medium difficulty, there is significant number of easy and hard puzzles. The number of reasoning steps follows a long-tail distribution, with many solutions requiring more than 5 steps and some hard puzzles requiring up to 30 steps of reasoning.
@@ 5.1 Experimental setup
We evaluate frontier closed-source reasoning models on PuzzleWorld, including GPT‑o3 (OpenAI, 2025), GPT‑4o (Achiam et al., 2023), Claude Opus 4 (Anthropic, 2025), Gemini-2.5-Pro (Comanici et al., 2025), Gemini-3-Pro (Google, 2025), and Grok 4 (xAI, 2025). We also evaluate open-source models Qwen QVQ (Qwen, 2024), InternVL3 (Zhu et al., 2025), and Kimi VL A3B (Team et al., 2025). We prompt each model with a comprehensive prompt as in Wang et al. (2025), followed by the puzzle images and transcribed flavor text. See Appendix F for the evaluation prompt.
We also provide a human baseline on PuzzleWorld, considering three tiers of puzzlehunter expertise: Novice, with no prior puzzlehunt experience; Enthusiasts, who showed interest or have occasionally participated (1-2 sessions) in puzzlehunts, and Experts among the top teams at monthly Puzzled Pint meetings. We we gathered 9 Novices and 9 Enthusiasts across high school and college ages. We sampled 5% puzzles from PuzzleWorldand assigned each participant to solve four puzzles. Participants were given an hour to solve each puzzle, matching the usual expected time at a live session, and were asked to provide paragraph explanations for their solution. For Experts, we use statistics from Puzzled Pint sessions in Syracuse, New York, and Bangalore, India, dating from January 2023 to June 2025. The statistics suggest that expert puzzlehunters consistently solve all five puzzles within one to two hours, which is on average less than the time prescribed to our human participants. We thus assume that human Experts achieve perfect accuracy on PuzzleWorld.
@@ 5.1.1 Automatic evaluation metrics
Beyond final answer accuracy, we additionally evaluate the models’ stepwise accuracy by comparing their solution with the annotated ground truth reasoning steps. Since puzzles can have multiple solution pathways, we define the stepwise accuracy score of a candidate solution to be the last annotated reasoning step it successfully executed out of all the reasoning steps. We implement an LLM judge (Zheng et al., 2023) with GPT‑4o to determine the stepwise score of each candidate solution. For each reasoning step in the reference solution, the LLM judge determines if the step is met by the candidate response. To evaluate LLM judge’s reliability, we compared its stepwise evaluations on 20 random puzzles against human evaluations. The LLM judge achieved a Pearson correlation of $r=0.829$ ($p=6.3\times 10^{-6}$) and a mean absolute error (MAE) of $0.083$ with respect to human scores, indicating strong alignment with human judgment.
@@ 5.2.1 Overall performance of frontier models
Table 2: Model performance. Accuracy (Acc) and stepwise accuracy (Step) are reported overall and per modality. Models struggle significantly on PuzzleWorld, most achieve only 1-4% answer accuracy, with closed-source models generally outperforming open-source ones. The best model, Gemini 3 Pro, solves only 18% of puzzles and reaches 40% stepwise accuracy, matching human Novice performance but significantly falling behind Enthusiasts.
@@
Claude Opus 4 4.50 24.56 4.20 23.77 4.04 22.60 4.37 26.93
Gemini 2.5 Pro 7.65 31.61 8.07 31.09 4.99 29.06 6.71 32.34
Gemini 3 Pro 18.00 39.99 20.30 39.34 14.71 38.81 20.25 39.99
Grok 4 3.33 13.79 3.85 13.64 3.70 14.19 1.56 11.22
Human Human Novice 13.89 23.10 16.98 25.32 11.00 22.70 16.67 24.92
Human Enthusiast 44.44 51.70 44.14 52.58 44.00 52.20 54.17 57.81
Human Expert 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
We report the models’ performance in Table 2. All models exhibit extremely low final answer accuracy on PuzzleWorld, with most achieving close to 1-4%. Gemini 3 Pro attains the highest overall accuracy at 18.00%, matching human Novice performance, while the best-performing open-source model, QVQ-72B-Preview, reaches just 1.36%. All models perform significantly worse than human Enthusiasts and Experts. Although the uniformly low accuracies underscore our benchmark’s difficulty, it offers limited insight into the models’ reasoning capabilities.
To address this, our stepwise evaluation metrics provide a more nuanced view of models’ reasoning performance. These metrics reveal that models with poor final answer accuracy, such as InternVL3, still demonstrate good intermediate reasoning, achieving up to 15.49% stepwise accuracy. Similarly, while QVQ-72B-Preview lags behind all closed-source models in final answer accuracy, it outperforms many of them in stepwise accuracy (30.2%), reflecting coherent and promising reasoning capabilities despite not reaching the correct final output. The combination of final answer and stepwise metrics enable PuzzleWorld to remain highly challenging while offering detailed diagnostics for model evaluation and development.
In terms of input modalities, models generally perform best on text-based puzzles, with significantly lower accuracy on puzzles involving unstructured visual inputs. Interestingly, most models achieve better performance on structured puzzles, such as crosswords where the spatial format constrained, over unstructured visual puzzles. In contrast, puzzles involving free-form visuals remain difficult, with models often achieving less than half their text puzzle accuracy on these inputs. Additionally, in Appendix D.2, we find that these modality-level performance discrepancies are not systematically driven by optical character recognition (OCR) limitations. These trends highlight current models’ persistent weaknesses in visual grounding and spatial reasoning. In contrast, we observe that visual and structured modalities do not pose additional difficulty for human solvers regardless of their puzzle solving experience.
We further analyze model performance by puzzle difficulty in Appendix E.2. We find that, as expected, model accuracy decreases as difficulty increases. Interestingly, GPT-o3 performs best on easy, structured puzzles – likely because of their clean, regular layouts – whereas even harder structured puzzles remain challenging as they become more irregular and visually complex.
5.3 Improving reasoning on downstream tasks with PuzzleWorld
To explore whether PuzzleWorld can support improvement in model reasoning, we fine-tuned an 8B InternVL3 model with supervised fine-tuning on annotated reasoning traces from 80% of the dataset, and evaluated performance on the 20% held-out test set. As a control, we fine-tuned the same model using only the final answers, without access to reasoning traces. Full training details are provided in Appendix F.2.
Table 3: Fine-tuning on PuzzleWorld. Fine-tuning InternVL3 on PuzzleWorld reasoning traces significantly improves its stepwise accuracy, while fine-tuning on final answer only causes model reasoning to collapse.
@@ 5.3 Improving reasoning on downstream tasks with PuzzleWorld
Our results in Table LABEL:tab:sft highlight the value of PuzzleWorld’s stepwise annotations. Fine-tuning on reasoning traces doubles the model’s stepwise accuracy – from the base model’s 4.78% to 11.00%. In contrast, fine-tuning on final answers alone impairs performance, reducing stepwise accuracy to 2.96% and driving answer accuracy to zero. Despite the improved stepwise accuracy, the fine-tuned model’s answer accuracy remained at 0.76%. This suggests that while fine-tuning enhanced model’s intermediate reasoning, it was insufficient to solve additional puzzles completely. This result underscores both the difficulty of PuzzleWorld and the limitations of straightforward fine-tuning approaches in addressing such complex reasoning challenges in our dataset.
Table 4: PuzzleWorld fine-tuned model performance on downstream reasoning tasks. Fine-tuning InternVL3 on PuzzleWorld leads to performance gains on visually-oriented tasks, such as Rebus puzzles, geometry, and visual question answering, but slightly reducing performance on problems more dependent on external knowledge, such as textbook question answering.
Dataset Task Base Model Fine-tuned on PuzzleWorld
Rebus Puzzles Puzzle reasoning 3.2% 5.1%
MathVista Geometry problem solving 65.87% 66.35%
Textbook question answering 63.92% 60.13%
Math word problem 62.37% 59.14%
Visual question answering 32.40% 39.11%
We then explore whether PuzzleWorld’s detailed stepwise annotation can improve models on downstream reasoning tasks. We finetuned a model on 80% of PuzzleWorld and evaluated it on two benchmarks: a Rebus puzzles dataset (Lee et al., 2025) involving visual metaphors without explicit instructions, and the MathVista dataset (Lu et al., 2024), ranging from general visual question answering to domain-specific geometry questions. Our results are shown in Table 4.
Finetuning on PuzzleWorld yields notable performance gains on Rebus puzzles, where the model’s accuracy increased from 3.2% to 5.1%. On MathVista, the fine-tuned model shows significant improvement in geometry problem solving and visual question answering, but its performance slightly decreased on tasks that are outside of PuzzleWorld’s reasoning skills and require external knowledge, such as textbook question answering and math word problems. This performance improvement suggests that the skills learned from PuzzleWorld are not merely task-specific. They represent transferable, general-purpose reasoning capabilities, making our dataset a reflection of diverse reasoning expertises and a valuable tool for enhancing models’ capabilities.
@@ 5.4 Detailed error analysis
Figure 6: Example puzzle errors. Left: (myopic reasoning) The model is unable to backtrack when it hits a dead end. Middle: (language bottleneck) The model misrepresents the visual contents due to inherent limitations of texts. Right: (sketching errors) The model fails to execute the visual sketching steps to obtain correct intermediate outputs. High-resolution images are in Appendix E.3.
In the following section, we highlight the main sources of errors by frontier reasoning multimodal LLMs on PuzzleWorld: myopic reasoning, limitations of language, and lack of visual sketching capabilities. We focus our analysis on GPT‑o3, and we provide qualitative examples for each error category in Figure 6. In Appendix D.2, we provide a less common ”visual forgetting” error pattern, where model deviates from visual puzzle content during extended reasoning processes.
@@
Figure 7: Stepwise accuracy distribution of GPT‑o3. GPT‑o3 receives stepwise accuracy of 0 for most puzzles, highlighting the model’s myopic reasoning tendencies and its inability to backtrack after committing to an incorrect first step.
Myopic reasoning.
@@
Despite strong performance on conventional benchmarks, frontier models often exhibit myopic commitment in their reasoning. Rather than exploring alternatives or revisiting prior steps, models tend to fixate on early, surface-level hypotheses, resulting in reasoning trajectories that are locally coherent but globally misaligned with the puzzle. For example, in Figure 6, solving the puzzle requires interpreting musical notes using a mix of binary, Morse code, and flag semaphores. Instead, GPT‑o3 identifies a Morse code reference early on and rigidly adheres to it–even as contradictions arise–demonstrating a lack of backtracking and verification.
@@
To further examine this behavior, we analyze the stepwise accuracy distribution of GPT‑o3 (Figure 7). We find that, on most puzzles, the model receive a score of 0, meaning the model often fails to correctly identify even the first step of the solution. Once committed to an incorrect path, the model rarely recovers, highlighting its brittle reasoning and a lack of dynamic self-correction, especially when it cannot rely on external environments for verification.
@@
Figure 8: Limitations of text. An example failure case where GPT‑o3 fails to represent a complex structured puzzle into text.
Limitations of language.
Modern multimodal models rely heavily on language-based reasoning strategies, such as chain-of-thought and code generation. However, this dependence becomes a bottleneck in puzzles with complex visual structure. In Figure 6, the puzzle is composed of four interlocking loops arranged in a clover-like pattern. This layout is visually intuitive, but difficult to represent in text.
@@
While GPT‑o3 correctly solves the word clues, it fails to capture the layout when converting the puzzle into text, as shown in Figure 8. This ultimately leads the model to derive an incorrect answer. This example highlights a broader limitation: when faced with highly complex structured inputs, models that default to textual reasoning often lose critical spatial information. This inherent mismatch between visual intuition and language-centric inference poses a fundamental challenge to models, especially those that depend on textual or code-based reasoning chains.
Figure 9: Reasoning skills of failed steps. We annotated the bottleneck steps with their reasoning skills, and we identify spatial reasoning as a primary bottleneck.
@@
While frontier models have made notable progress in logical deduction and arithmetic reasoning, they consistently underperform on spatial tasks that require sketching, drawing, and manipulating visual structure, such as decoding based on spatial arrangements or tracing paths through grids and mazes. In Figure 6, the model correctly solves the individual clues in a grid-based puzzle but fails to trace the intended path, resulting in an incorrect final answer. Humans naturally rely on sketching or mental imagery to reason through such spatial challenges, using external or internal visualizations to keep track of evolving structure. The absence of such capabilities in current models reveals a critical gap: without the ability to sketch and update a persistent visual representation, models are prone to failure in tasks that depend on spatial coherence.
To understand the impact of sketching to model performance, we manually analyzed 30 puzzles where GPT‑o3 produced incorrect answers. For each failure, we annotated the reasoning step responsible for the error with its corresponding reasoning skill. As shown in Figure 9, we found that 53.33% of these bottleneck steps involved spatial reasoning or sketching-related capabilities. This highlights a gap in models’ ability to manipulate visual structure during inference. Incorporating sketch-like visual memory and reasoning (Wu et al., 2024; Hu et al., 2024; Chen et al., 2025b) may offer a promising direction toward more robust and spatially grounded reasoning AI.